OpenStudy (anonymous):

simplify the expression sin 2 theta divided by 1+cos 2 theta

OpenStudy (tkhunny):

Is that \(\sin^{2}(\theta)\;or\;\sin(2\theta)\)?? How about the cosine, while we're at it?